Set out below are the financial statements of Emcar plc. Income Statement for the year ended 31 December 2017 $'000 Turnover 2,553 Cost of sales (1,814) Gross profit 739 25 Other Income: discount received Distribution costs Administrative expenses Finance cost (125) (264) (75) 300 Profit before tax Income tax expense Profit for the period 140 160 Statement of Financial Position as at 31 December 2017 2016 000.$ 000.$ Assets Non-current assets Property, plant and equipment Intangible assets Investments 380 305 250 200 25 630 530 Current assets Inventory Trade Receivables 150 102 390 315 Other Receivables Cash in hand 50 2 1 592 418 1,222 948 Total assets ===
2017 2016 Capital and reserves Share capital ($1 ordinary share) Share premium account Revaluation reserve Retained Earnings 000.$ 200 000.$ 150 160 150 100 91 260 180 Long term liabilities Long term loan 170 Current liabilities Trade payables 127 119 85 98 110 Bank overdraft Taxation 120 1,222 948 The following information is available (i) The proceeds for the sale of non-current asset investments amounted to $30,000 Property, plant and equipment with an original cost of $85, 000 and a net (ii book value of $45,000 were sold for $32,000 during the year. (iii) The following information relates to property, plant and equipment: At 31 December 31.12.2017 31.12.2016 000.$ 720 000.$ 595 Cost Accumulated depreciation 340 290 380 305 (iv) 50,000 $1 ordinary shares were issued during the year at a premium of $0.20 per share Dividends of $ 80,000 were paid during the year. REQUIRED Prepare the Cash Flow Statement for the year ended 31.12.17 D...C.fc 50
Answer #1
Cash Flow Statement
Indirect Method
Cash Flow from Operating Activities
Profit for the year $                            160.00
Adjustments
Depreciation $                    90.00 =340+85-45-290
Finance Cost $                    75.00 75
Loss on Sale of Investment $                      8.00 =85-45-32
Changes in Current Assets / Current Liabilities
Increase in Inventories $                  (48.00) =102-150
Increase in Accounts Receivables $                  (75.00) =315-390
Increase in Other Receivables $                  (50.00) =0-50
Increase in Trade Payables $                      8.00 =127-119
Increase in Taxation $                    10.00 =120-110
Total Adjustments $                              18.00
Cash from Operating Activities $                                   178.00
Cash flow from Investing Activities
Paid for Purchase of Equipment $                         (201.00) =(595-85-720)+100-91
Increase in Intangible Assets $                            (50.00) =200-250
Sale of Non Current Investments $                              30.00 30
Sale of Equipmet $                              32.00 32
Net cash used In investing activities $                                 (189.00)
Cash flow from Financing Activities
Borrowed by Long term Loan $                            120.00 =170-50
Finance Cost $                            (75.00)
Issue of Shares $                              60.00 =50*1.2
Ordinary Dividend Paid $                            (80.00)
Net cash used in financing activities $                                      25.00
Increase in Cash $                                      14.00
Opening Balance of Cash & Equivalents $                                   (97.00) =1-98
Ending Balance of Cash & Equivalents $                                   (83.00) =2-85

Bank Overdraft has been considered as cash & cash equivalent.